2 tonnes of galena is to be reduced to fine powder by passing through a crusher and a grinder in succession, drawing powder from the same drive. Screen analysis of feed, product from crusher and product from grinder indicated surface areas of 5, 115 and 1000 m²/kg respectively. If the power required by the drive to run the crusher driver assembly is 20 KW and efficiency of crusher is 20%, find the efficiency of grinder. Rittinger's number for galena is 100 m²/kj


1
Expert's answer
2021-06-18T07:55:00-0400

20kW = ∆A (115/5)

x kW = ∆A (1000/115)


x = "\dfrac{1000}{115}×20×\dfrac{5}{115}"


x = 7.56kW

7.56W = 20%

"\therefore" 100% = 37.8kW


2 tonnes = 2000kg

"2000kg ×\dfrac{1000}{115}m²/kg= 17391m²"


Rate of grinder reduction = "{37.8\u00d7}{100} =3780m\u00b2\/s"


"\%E =\dfrac{3780}{17391} ×100\%= 21.74\%"